OpenStudy (avengedslipknot):

Which of the following formulas find the sum of interior angles a polygon with n number of sides? A; 360n B;180n-360 C; 180m D; 180-180 WILL FAN AND MEDAL // I think its D but im not to sure

zepdrix (zepdrix):

You think it's D....? So you think the interior angles of a polygon will always add up to 180-180 degrees? :o

OpenStudy (avengedslipknot):

Then C cause interior angle formula basically equals 180 right ??

zepdrix (zepdrix):

A triangle has 3 sides. The interior angles of a triangle add up to 180 degrees. \(\large\rm 180\cdot1\) We have one of these 180 degrees for a triangle. I want to rewrite this 1 in a clever way that involves the number of sides. \(\large\rm 180\cdot(3-2)\) 1 can be rewritten as 3-2, now we have a setup that involves the number of sides (3). How about a square? A square has two triangles in it. So a square has two of these 180 degrees for interior angles. \(\large\rm 180\cdot2\) Again, we want to rewrite this 2 as something related to the number of sides of a square. \(\large\rm 180\cdot(4-2)\) So if we generalize this, we see that we're always multiplying 180 by `two less than` the number of sides (n). \(\large\rm 180\cdot(n-2)\)

zepdrix (zepdrix):

180(n-2) corresponds to one of your options, you'll have to distribute the 180 to figure out which one.

OpenStudy (avengedslipknot):

A

OpenStudy (denonakavro):

180(n-2) would distribute to 180n - 180(2) 180n-360